Share of Stock

At the beginning of the year, you purchased a share of stock for $55. Over the year the dividends paid on the stock were $2.50 per share.
 
a. Calculate the return if the price of the stock at the end of the year is $50. (Negative amount should be indicated by a minus sign. Round your answer to 2 decimal places. (e.g., 32.16))
b. Calculate the return if the price of the stock at the end of the year is $60. (Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1

$$ \begin{aligned} &\begin{array}{|l|l|} \hline \text { Part a: } & \\ \hline \text { Return=(P1+D-PO)/PO } & -4.55 \% \\ \hline(50+2.5-55) / 55 & \\ \hline \end{array}\\ &\begin{array}{|l|} \hline \text { Price at the end of year 1 } \\ \hline \text { Purchase price } \\ \hline \text { Dividend } \\ \hline \end{array}\\ &\begin{array}{|l|l|} \hline \text { Part b: } & \\ \hline \text { Return=(P1+D-PO)/PO } & 13.64 \% \\ \hline(60+2.5-55) / 55 & \\ \hline \end{array} \end{aligned} $$

answered by: TheSnow